# Asynchronous Feedback Network for Perceptual Point Cloud Quality Assessment

### v1

@@ Abstract
Recent years have witnessed the success of the deep learning-based technique in research of no-reference point cloud quality assessment (NR-PCQA). For a more accurate quality prediction, many previous studies have attempted to capture global and local features in a bottom-up manner, but ignored the interaction and promotion between them. To solve this problem, we propose a novel asynchronous feedback network (AFNet). Motivated by human visual perception mechanisms, AFNet employs a dual-branch structure to deal with global and local features, simulating the left and right hemispheres of the human brain, and constructs a feedback module between them. Specifically, the input point clouds are first fed into a transformer-based global encoder to generate the attention maps that highlight these semantically rich regions, followed by being merged into the global feature. Then, we utilize the generated attention maps to perform dynamic convolution for different semantic regions and obtain the local feature. Finally, a coarse-to-fine strategy is adopted to merge the two features into the final quality score. We conduct comprehensive experiments on three datasets and achieve superior performance over the state-of-the-art approaches on all of these datasets. The code will be available at https://github.com/zhangyujie-1998/AFNet.
@@ I Introduction
In the past years, deep-learning-based PCQA methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] have shown remarkable performance on multiple benchmarks, which can be applied directly for original 3D data or for projected 2D images. To achieve a more robust prediction, some visual-related methods [12, 4, 3, 6] advertise learning discriminative features, where two types of feature are usually considered: global feature and local feature. Many previous studies [12, 4], whether based on 3D or 2D operations, have sought to extract the two features in a bottom-up manner, seeing Fig. 1 (a). Given a multi-layer neural network, features of shallow and deep layers are separately extracted as local and global feature due to the gradually increasing receptive field size. The two types of feature are then simply fused (e.g., concatenation, addition) to regress into the final quality score.
@@
To address these problems, we draw two main inspirations from human visual perception mechanisms: i) recent studies [15, 16] have highlighted the distinct visual processing of global and local feature in different areas of the brain. Specifically, the left and right hemispheres of the brain are specialized in processing global and local information, respectively. These findings suggest the presence of a dual-stream processing mechanism in our brain, challenging existing methods that seek to capture distinct features using a single branch. ii) Mounting evidence [17, 18, 19] has indicated that global feature can first be perceived and then guide the extraction of local feature through a feedback connection, which is an asynchronous process. For example, when presented with a human body point cloud or its projected images, individuals can initially identify various semantic regions (e.g., head, trunk, or background) and then perform fine-grained observation for different regions. Based on these insights, we expect to design a new feature learning strategy for NR-PCQA to overcome the shortcomings of previous studies.
In this paper, we propose a new NR-PCQA method, called Asynchronous Feedback Network (AFNet). Motivated by the human visual perception mechanisms, AFNet employs a dual-branch structure to deal with global and local feature (called the global branch and the local branch), and a feedback module is constructed between branches, seeing Fig. 1 (b). More specifically, the input distorted point cloud is first projected into three types of multi-view images: texture images, depth images, and occupancy images. For the global branch, the texture and depth images are fed into a transformer-based encoder to generate the attention maps that highlight these semantically rich regions. The attention maps are then merged into the global feature through an occupancy-weighted multi-view fusion, which can reduce the influence of useless background information.
@@
  • We propose a new NR-PCQA method named AFNet, which is consistent with human visual perception. Through the feedback module between the two branches, the global feature can effectively guide the extraction of local feature.
@@ II Related Works
Global-Local Feature Learning. Global-local feature learning is a prevalent strategy in quality assessment research, which is based on the fact that perceptual quality is determined by both high-level semantic and low-level details. Many previous studies [12, 4] separately extract the output of shallow and deep layers of networks as local and global feature and perform parallel processing, ignoring the interaction and promotion between features. In image quality assessment scenarios, researchers have proposed several methods to better utilize the two types of features. An intriguing way [26] is to utilize the global feature to generate a hyper network [27] to replace the common quality regression module, where the hyper network represents the perception rule for a specific sample. In [28] and [29], Chen et al. and Hu et al. both proposed cross-scale fusion mechanisms that allow information to propagate between features. However, the above improvements focus more on feature post-processing, that is, how to increase the interaction between already generated features, ignoring the intrinsic temporal order of feature extraction in brain mechanisms. In comparison, we intend to use global feature directly to guide the ”generation” of local feature, which is an asynchronous process and shows more conformity to visual perception.
Figure 2: The proposed AFNet framework, which includes two branches and a feedback module. We first divide images into patches and feed them into a transformer encoder, and the attention maps of class token are merged into the global feature. Then, we utilize the attention maps to generate a guided mask and multiple groups of filter, followed by performing region-aware convolution to derive the local feature. Finally, we employ a coarse-to-fine quality prediction based on the extracted two features.
@@ III-B Preprocessing
Instead of inferring point cloud quality directly in 3D space, we project point clouds into 2D images to better leverage mature 2D pre-trained models targeted for semantic-related tasks (e.g., classification). Given a point cloud $X$, we render it into multi-view images from six perpendicular perspectives (that is, along the positive and negative directions of the x,y,z axes). To reduce information loss during the projection process, we obtain three different types of image to represent $X$: texture images, depth images, and occupancy images. The former two types of image reflect the original color and geometry information and are normalized to the range $[0,1]$. The occupancy images are binary images locating the point cloud contents, where ”1” indicates the point cloud contents while ”0” indicates background areas. We separately denote the three types of multi-view images of $X$ as $\{X_{t}^{i}\in\mathbb{R}^{H\times W\times 3}|_{i=1}^{6}\}$, $\{X_{d}^{i}\in\mathbb{R}^{H\times W}|_{i=1}^{6}\}$ and $\{{X}_{o}^{i}\in\mathbb{R}^{H\times W}|_{i=1}^{6}\}$, where $i$ denotes the $i$-th viewpoint.
Figure 3: Visusalization of the attention maps corresponding to the class token and the generated mask. (a) Original projected texture images. (b) Attention maps corresponding to the class token. (c) Guided masks.
@@ III-C Attention-aware Global Feature Extraction
where ${x}\in\mathbb{R}^{D}$ is the token embedding; $\mathrm{PatchEmbed}_{t(d)}(\cdot)$ are two independent embedding operations consisting of a trainable linear projection after patch flattening.
Next, we follow [30] to prepend a learnable embedding ${x}_{c}$ named class token to the sequence of embedded patches. The token sequence can be represented as ${z}_{0}=[{x}_{c};{x}_{1};{x}_{2};\cdots;{x}_{N}]\in\mathbb{R}^{(N+1)\times D}$, where $N=HW/P^{2}$ denotes the patch number. We then input ${z}_{0}$ into the ViT, which consists of $L$ consecutive identical transformer blocks and is initialized with the optimized weights on ImageNet. Taking the $l$-th block as an example, it takes the output ${z}_{l-1}$ of the previous $(l-1)$-th block as input to compute its output ${z}_{l}$:
$\tilde{{z}}_{l}=\mathrm{MSA}(\mathrm{LN}({z}_{l-1}))+{z}_{l-1},$ (2)
${z}_{l}=\mathrm{MLP}(\mathrm{LN}(\tilde{{z}}_{l}))+\tilde{{z}}_{l},$
where $\mathrm{MSA(\cdot)}$ denotes the multi-head self-attention operation with $N_{h}$ parallel heads; $\mathrm{LN(\cdot)}$ denote the layer normalization [34]. $\mathrm{MLP}(\cdot)$ represents the multi-layer perceptron.
Attention Map Extraction. The ViT effectively models the relationship between patches using the self-attention operation [35]. For the $j$-th head of $\mathrm{MSA}(\cdot)$, the computational process can formulated as follows:
@@
where $\mathrm{SA_{j}(\cdot)}$ represent the $j$-th self-attention operation, $j\in[1,N_{h}]$; $\Gamma_{j}$ denotes the projection matrix; $A_{j}$ denotes the attention matrix. According to Eq. (3), we can see that each element of the attention matrix $A_{j}$ indicates the pairwise correlation between two patch representations in ${z}_{l-1}$. Furthermore, the class token collects information from all patches via the self-attention operation. In traditional classification tasks, the output of the class token (i.e., $z_{L}[:,1]$) will be transformed into a semantic prediction via a small MLP [30]. Therefore, the correlation between the class token and an arbitrary patch representation can effectively reflect the relevance between the patch and semantics, which facilitates the generation of needed global feature. We define the attention maps $A_{c}$ of the class token as:
@@
where $w_{i}$ denotes the occupancy ratio, that is, the proportion of ”1” in the occupancy image $X_{o}^{i}$. We take $f_{g}\in\mathbb{R}^{D_{o}}$ as the final global feature.
III-D Feedback-guided Local Feature Extraction
Specifically, the attention maps of different viewpoints are first stitched as $A_{s}\in\mathbb{R}^{\frac{2H}{P}\times\frac{3W}{P}\times N_{h}}$. Considering that the occupancy images directly distinguish the blank backgrounds using ”0”, it is reasonable to merge the occupancy information into the attention maps to achieve better region division. For this purpose, the stitched occupancy image $S_{o}$ is first resized to the size of $\frac{2H}{P}\times\frac{3W}{P}$, and then multiply with each channel of $A_{s}$, i.e.,
@@ III-E Coarse-to-fine Quality Prediction
According to Eq. (8), when the cosine similarity between $f_{g}$ and $f_{l}$ becomes 0, it indicates that there is no linear correlation between the two features.. A margin was given to the loss function because being exactly opposite could harm the representation ability [37].
@@
where the superscript ”$b$” denotes the $b$-th samples in a mini-batch with the size of $B$. To achieve a progressive quality prediction, we further propose a branch-wise quality ranking loss as:
@@ IV-A Databases and Evaluation Protocols
AFNet 0.957 0.930 0.678 0.897 0.899 9.971 0.690 0.680 0.583
Figure 6: Exemplary point clouds with the subjective MOS, the predicted quality score of different NR-PCQA metrics. The predictions that are closest to the MOS are highlighted in boldface.
@@ IV-B Implementation Details
IV-C Performance Comparison
Overall Performance Comparison. Table II lists the experimental results on the three PCQA databases in terms of SROCC, PLCC, and RMSE. The proposed method, AFNet, is compared with 11 FR-PCQA metrics, and 7 NR-PCQA metrics. we run all published codes to produce the results. We strictly retrain the available baselines with the same database split set up to keep the comparison fair. In addition, for the FR-PCQA methods that require no training, we simply validate them on the same testing sets and record the average performance.
TABLE III: Cross-dataset evaluation for NR-PCQA methods. Training and testing are both conducted on complete datasets. Results of PLCC are reported.
Train Test PQA-Net GPA-Net ResSCNN IT-PCQA MM-PCQA GMS-PCQA 3DTA AFNet
@@ IV-C Performance Comparison
For each database, the top two performance results for each evaluation criterion are highlighted in boldface and underline. We can see that the proposed AFNet exhibits outstanding performance across all three datasets. Especially, the proposed AFNet presents remarkable performance on the LS-PCQA, a database with wide diversity in terms of both content and distortion type, which demonstrates the adaptability and robustness of our model. In comparison, the performance of the existing FR and NR methods varies significantly across different datasets. For example, GMS-PCQA demonstrates strong performance on LS-PCQA but exhibits poor performance on WPC; 3DTA works well on SJTU-PCQA but presents inferior performance on LS-PCQA. IT-PCQA performs poorly because it is based on domain adaptation, in which the labels of point clouds are not utilized in the training. Meanwhile, it should be noted that our method presents competitive performance compared to the FR-PCQA methods despite the inaccessibility of the reference information.
Furthermore, we show some examples of distorted point clouds with the subjective MOS, the predicted score of the proposed AFNet and other NR-PCQA models in Fig. 6, where two point clouds are impaired by G-PCC trisoup compression and additive noise. It can be seen that the predicted score provided by AFNet aligns well with subjective perception. In comparison, other models are not sensitive to quality changes because their predicted scores for two point clouds are relatively close.
Qualitative Analysis. To demonstrate that our model can attentively capture the critical features related to distortions, we also present a qualitative analysis using a prevalent tool for network explanation, that is, Grad-CAM++ [50]. As illustrated in Fig. 7, subfigure (a) presents three samples with the same content but impaired by three different distortions (i.e., Rayleigh noise, V-PCC compression, and G-PCC compression), and subfigure (b) presents their Grad-CAM++ maps that reflect what makes the network perform decision. Note that the extracted activation maps correspond to the last convonlutional layer on the local branch (i.e., the $1\times 1$ Conv in Fig. 4 (b)). From the figure, we have the following observations: i) the activation map associated with Rayleigh noise distortion mostly highlights these flat regions. This is mainly because Rayleigh noise is located at isolated points. According to texture masking effect [51], noise in complex texture regions (e.g., orange box) is generally harder to perceive than noise in homogeneous regions (e.g. blue box). ii) For two types of compression distortion, attribute quantization in V-PCC leads to abnormal color mutations, whereas Octree decomposition in G-PCC causes a sparse point distribution. Consequently, the activation maps associated with V-PCC and G-PCC emphasize the whole point cloud because these deformations are severe and not localized. Based on these instances, we can see that the proposed AFNet can effectively capture and distinguish various distortions.
@@
TABLE IV: Ablation study for the key modules on the WPC dataset. ‘✓’ or ‘✗’ means the setting is preserved or discarded.
Index Global Local Feedback PLCC SROCC RMSE
(1) 0.8291 0.8233 12.6480
(2) 0.3940 0.2766 20.8121
(3) 0.8701 0.8692 11.1941
(4) 0.8974 0.8987 9.9707

### v2

@@ Abstract
Recent years have witnessed the success of the deep learning-based technique in research of no-reference point cloud quality assessment (NR-PCQA). For a more accurate quality prediction, many previous studies have attempted to capture global and local features in a bottom-up manner, but ignored the interaction and promotion between them. To solve this problem, we propose a novel asynchronous feedback quality prediction network (AFQ-Net). Motivated by human visual perception mechanisms, AFQ-Net employs a dual-branch structure to deal with global and local features, simulating the left and right hemispheres of the human brain, and constructs a feedback module between them. Specifically, the input point clouds are first fed into a transformer-based global encoder to generate the attention maps that highlight these semantically rich regions, followed by being merged into the global feature. Then, we utilize the generated attention maps to perform dynamic convolution for different semantic regions and obtain the local feature. Finally, a coarse-to-fine strategy is adopted to merge the two features into the final quality score. We conduct comprehensive experiments on three datasets and achieve superior performance over the state-of-the-art approaches on all of these datasets. The code will be available at https://github.com/zhangyujie-1998/AFQ-Net.
@@ I Introduction
In the past years, deep-learning-based PCQA methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] have shown remarkable performance on multiple benchmarks, which can be applied directly for original 3D data or for projected 2D images. To achieve a more robust prediction, some visual-related methods [13, 4, 3, 6, 14] advertise learning discriminative features, where two types of feature are usually considered: global feature and local feature. Many previous studies [13, 4], whether based on 3D or 2D operations, have sought to extract the two features in a bottom-up manner, seeing Fig. 1 (a). Given a multi-layer neural network, features of shallow and deep layers are separately extracted as local and global feature due to the gradually increasing receptive field size. The two types of feature are then simply fused (e.g., concatenation, addition) to regress into the final quality score.
@@
To address these problems, we draw two main inspirations from human visual perception mechanisms: i) recent studies [17, 18] have highlighted the distinct visual processing of global and local feature in different areas of the brain. Specifically, the left and right hemispheres of the brain are specialized in processing global and local information, respectively. These findings suggest the presence of a dual-stream processing mechanism in our brain, challenging existing methods that seek to capture distinct features using a single branch. ii) Mounting evidence [19, 20, 21] has indicated that there exists an asynchronous process in visual perception. In general, human visual cortex can first perceive the global feature, which is then used to guide the extraction of the local feature through a feedback connection. For example, when presented with a human body point cloud or its projected images, individuals can initially identify various semantic regions (e.g., head, trunk, or background) and then perform fine-grained observation for different regions. Based on these insights, we expect to design a new feature learning strategy for NR-PCQA to overcome the shortcomings of previous studies.
In this paper, we propose a new NR-PCQA method, called Asynchronous Feedback Quality Prediction Network (AFQ-Net). Motivated by the human visual perception mechanisms, AFQ-Net employs a dual-branch structure to deal with global and local feature (called the global branch and the local branch), and a feedback module is constructed between branches, seeing Fig. 1 (b). More specifically, the input distorted point cloud is first projected into three types of multi-view images: texture images, depth images, and occupancy images. For the global branch, the texture and depth images are fed into a transformer-based encoder to generate the attention maps that highlight these semantically rich regions. The attention maps are then merged into the global feature through an occupancy-weighted multi-view fusion, which can reduce the influence of useless background information.
@@
  • We propose a new NR-PCQA method named AFQ-Net, which is consistent with human visual perception. Through the feedback module between the two branches, the global feature can effectively guide the extraction of local feature.
@@ II Related Works
Recently, some research [28, 29, 14, 30, 31, 32] has aimed to measure point cloud quality using payload information extracted from the compressed bitstream, which benefits real-time and nonintrusive quality monitoring. For example, Liu et al. [29] studied the relationship between the perceptual quality and the texture/geometry quantization parameter in video-based point cloud compression (V-PCC) algorithm and proposed a no-reference bitstream-layer model. In [30], Lv et al. proposed a new texture distortion model for lossless geometric coding in geometry-based point cloud compression (G-PCC) and integrated it with the position quantization scale (PQS) results for quality prediction.
Figure 2: The proposed AFQ-Net framework, which includes two branches and a feedback module, where $\oplus$ denotes the concatenation operation. We first divide images into patches and feed them into a transformer encoder, and the attention maps of class token are merged into the global feature. Then, we utilize the attention maps to generate a guided mask and multiple groups of filter, followed by performing region-aware convolution to derive the local feature. Finally, we employ a coarse-to-fine quality prediction based on the extracted two features.
Global-Local Feature Learning. Global-local feature learning is a prevalent strategy in quality assessment research, which is based on the fact that perceptual quality is determined by both high-level semantic and low-level details. Many previous studies [13, 4] separately extract the output of shallow and deep layers of networks as local and global feature and perform parallel processing, ignoring the interaction and promotion between features. In image/video quality assessment scenarios, researchers have proposed several methods to better utilize the two types of feature. An intriguing way [33] is to utilize the global feature to generate a hyper network [34] to replace the common quality regression module, where the hyper network represents the perception rule for a specific sample. In [35] and [36], Chen et al. and Hu et al. both proposed cross-scale fusion mechanisms that allow information to propagate between features. In [37], Chen et al. proposed an attention module to assign weight for multi-scale features by their importance. However, the above improvements focus more on feature post-processing, that is, how to increase the interaction between already generated features, ignoring the intrinsic temporal order of feature extraction in brain mechanisms. In comparison, we intend to use global feature directly to guide the “generation” of local feature, which is an asynchronous process and shows more conformity to visual perception.
@@ III-A Overview
Figure 3: Visusalization of the attention maps corresponding to the class token and the generated mask. (a) Original projected texture images. (b) Attention maps corresponding to the class token. (c) Guided masks that divide the images into multiple semantic regions. Different colors in (b) indicate the size of the element value in the attention map; the brighter the color, the larger the element value. In (c), colors are used to distinguish different regions, and pixels of the same color are considered to have the same semantics.
@@ III-B Preprocessing
Instead of inferring point cloud quality directly in 3D space, we project point clouds into 2D images to better leverage mature 2D pre-trained models targeted for semantic-related tasks (e.g., classification). Given a point cloud $X$, we render it into multi-view images from six perpendicular perspectives (that is, along the positive and negative directions of the x,y,z axes). To reduce information loss during the projection process, we obtain three different types of image to represent $X$: texture images, depth images, and occupancy images. More specifically, we use the PointsRenderer function in the Pytorch3D [40] library for point cloud projection. The renderer uses an orthographic camera, and applies alpha compositing to generate the pixel values (please refer to [41] for more details). The acquired former two types of image reflect the original color and geometry information and are normalized to the range $[0,1]$. The occupancy images are binary images locating the point cloud contents, where “1” indicates the point cloud contents while “0” indicates background areas. We separately denote the three types of multi-view images of $X$ as $\{X_{t}^{i}\in\mathbb{R}^{H\times W\times 3}|_{i=1}^{6}\}$, $\{X_{d}^{i}\in\mathbb{R}^{H\times W}|_{i=1}^{6}\}$ and $\{{X}_{o}^{i}\in\mathbb{R}^{H\times W}|_{i=1}^{6}\}$, where $i$ denotes the $i$-th viewpoint.
@@ III-C Attention-aware Global Feature Extraction
where ${x}\in\mathbb{R}^{D}$ is the token embedding; $\mathrm{PatchEmbed}_{t(d)}(\cdot)$ are two independent embedding operations.
Next, we follow [38] to prepend a learnable embedding ${x}_{c}$ named class token to the sequence of embedded patches. The token sequence can be represented as ${z}_{0}=[{x}_{c};{x}_{1};{x}_{2};\cdots;{x}_{N}]\in\mathbb{R}^{(N+1)\times D}$, where $N=HW/P^{2}$ denotes the patch number. We then input ${z}_{0}$ into the ViT, which consists of $L$ consecutive identical transformer blocks and is initialized with the optimized weights on ImageNet. Taking the $l$-th block as an example, it takes the output ${z}_{l-1}$ of the previous $(l-1)$-th block as input to compute its output ${z}_{l}$.
Attention Map Extraction. The ViT effectively models the relationship between patches using multi-head self-attention operation $\mathrm{MSA}(\cdot)$ [44]. For the $j$-th head of $\mathrm{MSA}(\cdot)$, the computational process can formulated as follows:
@@
where $\mathrm{SA_{j}(\cdot)}$ represent the $j$-th self-attention operation, $j\in[1,N_{h}]$; $\mathrm{LN(\cdot)}$ denote the layer normalization [45]; $\Gamma_{j}$ denotes the projection matrix; $A_{j}$ denotes the attention matrix. According to Eq. (2), we can see that each element of the attention matrix $A_{j}$ indicates the pairwise correlation between two patch representations in ${z}_{l-1}$. Furthermore, the class token collects information from all patches via the self-attention operation. In traditional classification tasks, the output of the class token (i.e., $z_{L}[:,1]$) will be transformed into a semantic prediction via a small MLP [38]. Therefore, the correlation between the class token and an arbitrary patch representation can effectively reflect the relevance between the patch and semantics, which facilitates the generation of needed global feature. We define the attention maps $A_{c}$ of the class token as:
@@
where $w_{i}$ denotes the occupancy ratio, that is, the proportion of “1” in the occupancy image $X_{o}^{i}$. We take $f_{g}\in\mathbb{R}^{D_{o}}$ as the final global feature.
@@
III-D Feedback-guided Local Feature Extraction
Specifically, the attention maps of different viewpoints are first stitched as $A_{s}\in\mathbb{R}^{\frac{2H}{P}\times\frac{3W}{P}\times N_{h}}$. Considering that the occupancy images directly distinguish the blank backgrounds using “0”, it is reasonable to merge the occupancy information into the attention maps to achieve better region division. For this purpose, the stitched occupancy image $S_{o}$ is first resized to the size of $\frac{2H}{P}\times\frac{3W}{P}$, and then multiply with each channel of $A_{s}$, i.e.,
@@ III-E Coarse-to-fine Quality Prediction
According to Eq. (7), when the cosine similarity between $f_{g}$ and $f_{l}$ becomes 0, it indicates that there is no linear correlation between the two features. A margin was given to the loss function because being exactly opposite could harm the representation ability [47].
@@
where the superscript “$b$” denotes the $b$-th samples in a mini-batch with the size of $B$. To achieve a progressive quality prediction, we further propose a branch-wise quality ranking loss as:
@@ IV-A Databases and Evaluation Protocols
AFQ-Net 0.957 0.930 0.678 0.897 0.899 9.971 0.690 0.680 0.583
@@ IV-B Implementation Details
TABLE III: PERFORMANCE COMPARISON FOR NR-PCQA METRICS ON M-PCCD and BASICS databases
M-PCCD PQA-Net GPA-Net ResSCNN IT-PCQA MM-PCQA GMS-3DQA 3DTA AFQ-Net
PLCC 0.554 0.383 0.298 0.310 0.829 0.818 0.703 0.816
SROCC 0.478 0.396 0.267 0.315 0.810 0.826 0.765 0.834
RMSE 0.886 0.847 1.004 0.996 0.693 0.702 0.995 0.740
BASICS PQA-Net GPA-Net ResSCNN IT-PCQA MM-PCQA GMS-3DQA 3DTA AFQ-Net
PLCC 0.655 0.391 0.367 0.465 0.793 0.895 0.896 0.905
SROCC 0.387 0.352 0.344 0.369 0.738 0.807 0.825 0.793
RMSE 0.797 0.975 0.997 0.878 0.628 0.472 0.488 0.449
Figure 6: Exemplary point clouds with the subjective MOS and the predicted quality score of different NR-PCQA metrics. The predictions that are closest to the MOS are highlighted in boldface.
IV-C Performance Comparison
Overall Performance Comparison. Table II lists the experimental results on META-3D . The proposed method, AFQ-Net, is compared with 11 FR-PCQA metrics, and 7 NR-PCQA metrics. we run all published codes to produce the results. We strictly retrain the available baselines with the same database split set up to keep the comparison fair. In addition, for the FR-PCQA methods that require no training, we simply validate them on the same testing sets and record the average performance.
@@ IV-C Performance Comparison
For each database, the top two performance results for each evaluation criterion are highlighted in boldface and underline. We can see that the proposed AFQ-Net exhibits outstanding performance across all three datasets. Especially, the proposed AFQ-Net presents remarkable performance on the LS-PCQA, a database with wide diversity in terms of both content and distortion type, which demonstrates the adaptability and robustness of our model. In comparison, the performance of the existing FR and NR methods varies significantly across different datasets. For example, GMS-3DQA demonstrates strong performance on LS-PCQA but exhibits poor performance on WPC; 3DTA works well on SJTU-PCQA but presents inferior performance on LS-PCQA. IT-PCQA performs poorly because it is based on domain adaptation, in which the labels of point clouds are not utilized in the training. Meanwhile, it should be noted that our method presents competitive performance compared to the FR-PCQA methods despite the inaccessibility of the reference information.
To further validate the model performance in compression scenarios, we test the model performance on two PCQA databases that consist only of compressed samples, i.e., the M-PCCD database (232 distorted samples corresponding to 8 references) [59] and the BASICS database (1494 distorted samples corresponding to 75 references) [60]. 4-fold and 5-fold cross-validations are applied for the two databases, respectively. We report the performance of different NR-PCQA metrics on M-PCCD and BASICS in Tab III. According to the table, we can see that AFQ-Net provides the best SROCC on M-PCCD and the best PLCC and RMSE on BASICS, which demonstrates that the proposed method is sensitive to compression distortions. MM-PCQA and 3DTA both presents inferior performance on one database (MM-PCQA on BASICS and 3DTA on M-PCCD). GMS-3DQA works well, which is also based on 2D projection. Based on the results, it seems that metrics using 3D information present unstable performance on different databases. This may be because those metrics only utilize some local patches for quality prediction, which lacks global perception and is easily affected by point density. However, for those metrics that only leverage 2D projection, how to mimic different rendering strategies for different databases has not yet been explored in depth.
We also show some examples of distorted point clouds with the subjective MOS, the predicted score of the proposed AFQ-Net and other NR-PCQA models in Fig. 6, where two point clouds are impaired by G-PCC trisoup compression and additive noise. It can be seen that the predicted score provided by AFQ-Net aligns well with subjective perception. In comparison, other models are not sensitive to quality changes because their predicted scores for two point clouds are relatively close.
Qualitative Analysis. To demonstrate that our model can attentively capture the critical features related to distortions, we also present a qualitative analysis using a prevalent tool for network explanation, that is, Grad-CAM++ [61]. As illustrated in Fig. 7, subfigure (a) presents three samples with the same content but impaired by three different distortions (i.e., Rayleigh noise, V-PCC compression, and G-PCC compression), and subfigure (b) presents their Grad-CAM++ maps that reflect what makes the network perform decision. Note that the extracted activation maps correspond to the last convonlutional layer on the local branch (i.e., the $1\times 1$ Conv in Fig. 4 (b)). From the figure, we have the following observations: i) the activation map associated with Rayleigh noise distortion mostly highlights these flat regions. This is mainly because Rayleigh noise is located at isolated points. According to texture masking effect [62], noise in complex texture regions (e.g., orange box) is generally harder to perceive than noise in homogeneous regions (e.g., blue box). ii) For two types of compression distortion, attribute quantization in V-PCC leads to abnormal color mutations, whereas Octree decomposition in G-PCC causes a sparse point distribution. Consequently, the activation maps associated with V-PCC and G-PCC emphasize the whole point cloud because these deformations are severe and not localized. Based on these instances, we can see that the proposed AFQ-Net can effectively capture and distinguish various distortions.
TABLE IV: Cross-dataset evaluation for NR-PCQA methods. Training and testing are both conducted on complete datasets. Results of PLCC are reported.
Train Test PQA-Net GPA-Net ResSCNN IT-PCQA MM-PCQA GMS-3DQA 3DTA AFQ-Net
@@
WPC SJTU 0.255 0.549 0.568 0.395 0.740 0.517 0.726 0.558
SJTU WPC 0.266 0.420 0.271 0.310 0.301 0.523 0.339 0.385
Cross-Dataset Evaluation. The cross-dataset evaluation is conducted to test the generalizability of the NR-PCQA methods when encountering various data distributions. In Table IV, we first train the compared models on one database and test the trained model on another database, and then swap the training set and the test set. Note that we use the whole database for training or testing. For each validation, we record the result with minimal training loss and report the results in Table IV. From the table, we can see that: i) AFQ-Net achieves good performance when trained on LS-PCQA, which shows the generalizability of our model when rich content categories are provided. ii) Almost all methods (including AFQ-Net) show poor performance when trained on SJTU-PCQA and tested on LS-PCQA and WPC (denoted by SJTU2LS and SJTU2WPC validation). This is largely due to the limited and extremely biased content in SJTU-PCQA (5 out of 9 contents are about the human body). iii) AFQ-Net trained on WPC shows opposite trends when validated on SJTU-PCQA and LS-PCQA. More specifically, AFQ-Net performs the best when validated on LS-PCQA but presents relatively inferior performance when validated on SJTU-PCQA. This may be due to the large difference in content categories between SJTU-PCQA and WPC (human body vs. object). Notably, some recent methods demonstrate superior cross-dataset performance on several validations. For instance, MM-PCQA and 3DTA perform well on the WPC2SJTU validation, likely due to their use of 3D information as input, which enables them to capture common density variations in SJTU-PCQA more effectively. GMS-3DQA achieves the best results on the SJTU2WPC validation, potentially because of its mini-patch sampling strategy. The mini-patch map in GMS-3DQA preserves the distortion patterns of point clouds while blurring content information, which appears advantageous when there are substantial content differences between databases. To enhance the performance of AFQ-Net, a promising approach would be to replace the stitched image in the local branch with the mini-patch map, thus improving its capacity to evaluate local distortions. Additionally, directly applying the asynchronous feedback strategy for 3D data is also a practical and intriguing avenue.
@@
TABLE V: Ablation study for the key modules. ‘✓’ or ‘✗’ means the setting is preserved or discarded.
Index Global Local Feedback WPC LS-PCQA
PLCC SROCC PLCC SROCC
(1) 0.8291 0.8233 0.6736 0.6512
(2) 0.3940 0.2766 0.4502 0.4129
(3) 0.8701 0.8692 0.6786 0.6662
(4) 0.8974 0.8987 0.6902 0.6796
IV-D Ablation Studies
Ablation Study for the Key Modules. To verify the effectiveness of the key modules in the AFQ-Net, we further investigate the contribution of different components and report the results in Table V. In the table, the index (1) and (2) denote using only global branch or local branch for quality prediction respectively. (3) indicates one dual-branch network without feedback connection (i.e., replacing the DRConv in Fig. 4 (b) with $3\times 3$ convolution), while (4) represents the full model. From the table, we have the following observations: i) Seeing (1) and (2), we see that using only the global branch performs better than using only the local branch structure. This is mainly because the parameter size of the global branch is significantly larger than that of the local branch; consequently, the global branch has stronger fitting ability. Meanwhile, the pre-trainning on ImageNet inject the global branch some quality-related prior knowledge, which boosts its performance. ii) Comparing (1) or (2) with (3), the dual-branch structure performs significantly better than the single-branch structure, demonstrating the complementarity of the two features; iii) Seeing (3) and (4), the feedback module further brings a noticeable gain, which verifies the effectiveness of the feedback mechanism in visual quality assessment.